# The Aharony-Bergman-Jafferis-Maldacena theory on a circle

### v5

@@ Abstract
In this work, we bootstrap the 4-point correlators on the 1D celestial circle using 3D symmetries in the Aharony-Bergman-Jafferis-Maldacena theory as constraints. We find that the dual inversion property is strong enough to replace the crossing symmetry condition (or cyclic invariant condition) when bootstrapping. We also give some results about the conformal block expansion coefficients which contain the spectrum and the leading multi-OPE of this celestial CFT. Furthermore, we find a non-local 2-particle operator from the multi-OPE at the leading OPE. Although we studied a specific theory, the methods used are valid for general cases.
@@ 1 Introduction and review
Recently, celestial holography, which translates the 4D amplitudes in the bulk to CFT correlators on the celestial sphere that is equivalent to the lightlike infinity of the origin [12, 13], has given us many illustrating results [14, 15, 16, 17, 18]. Inspired by AdS/CFT, celestial holography aims to help us understand quantum gravity. In addition, it also helps us understand other theories from a new aspect. For massless amplitudes, this corresponding can be realized by the Mellin transformation on the magnitude of the external momenta, namely (for 4D amplitudes)
@@
One may ask if we can obtain more information on the ABJM from a celestial perspective. Furthermore, can we bootstrap the celestial CFT (CCFT) corresponding to ABJM? In this paper, we will answer these two questions tentatively. We will first show how to bootstrap the 4-point correlators of the CCFT using the 3D symmetry as constraints and then show the conformal block expansion of the 4-point correlator and the loop correction. Finally, we will show the multi-collinear behavior [20, 21] and the in-in (or out-out) OPE of this CCFT.
This paper is organized as follows. In section 2, we will use the celestial version of 3D spin-helicity formalism, which was proposed in [22], to express the 3D symmetries as 1D constraints for tree-level celestial amplitudes [^1] As a supplement, for the celestial version of 4D massive spinor-helicity formalism, see [23].. We will also write down the ansatz for the 4-point tree-level celestial amplitudes. In section 3, we will solve these constraints. In section 4, we will first introduce how the dual inversion acts on the celestial variables for 4-point amplitudes. In section 5, we will investigate the spectrum of this CCFT and the influence of the loop corrections. In section 6, we will consider the multi-collinear limit to obtain the celestial OPE.
@@
$P^{ab}\mathcal{A}_{n}$ $=\sum_{i}\epsilon_{i}x_{i}^{j}\mathcal{T}_{i}\mathcal{A}_{n}=0,\text{ for $j=0,1,2$.}$ (2.4)
@@
$Q^{aA}\mathcal{A}_{n}$ $=-\sum_{i}\sqrt{-2\epsilon_{i}}\mathcal{T}_{i}^{1/2}\left(\begin{array}[]{c}1\\ x_{i}\end{array}\right)\partial_{\eta_{i}^{A}}\mathcal{A}_{n}=0.$
where the operator $\mathcal{T}_{i}$ can add 1 to the $\Delta_{i}$ in the 1D correlator $\mathcal{A}_{n}$. In ABJM, besides the usual superconformal symmetry, one can also find the dual superconformal symmetry [10] which will be manifest after using the dual coordinates. Such dual superconformal symmetry can generate a larger algebra together with the original superconformal symmetry, which is a Yangian [24]. This fact is similar to the case in the $\mathcal{N}=4$ SYM, and we only need to consider two level-1 operators in the Yangian. When acting on the 1D correlators, they are equivalent to the dual conformal boost operator and its superpartner. The dual conformal boost operator for ABJM amplitudes reads [10]
@@
    It has the form
    $\mathcal{A}_{4}=g(\epsilon_{i},\Delta_{i})f(z,\eta_{i}^{A})\prod_{i<j}x_{ij}^{\Delta/3-\Delta_{i}-\Delta_{j}}$
    with $z=\frac{x_{12}x_{34}}{x_{13}x_{24}}$ the cross-ratio and $\Delta=\sum_{i=1}^{4}\Delta_{i}$.
@@ 3 Four-point correlators from 1D constraints
[…]
@@ 4 Dual inversion
up to a constant to obtain (4.4). Hence for a 1D CFT, we can impose the constraints (4.4) and also define the dual inversion (LABEL:dual2) and (4.10) for $x_{ij}$ and $\eta_{i}^{A}$ respectively. This constraint, together with the first 3 ansatz in section 2, will determine the 4-point correlator of the 1D CCFT dual to the 3D ABJM theory. Note that the properties (LABEL:dual2) and (4.10) are valid for 4-point correlators in any 1D CFT.
@@ 5 Conformal block expansion and loop corrections
Here $\mu$ is the regularization scale and $\zeta(2)=\pi^{2}/6$ is the Riemann zeta function $\zeta(z)$. After acting this operator on the 4-point correlator $\mathcal{A}_{4}$ we will obtain the loop correction. Such a logarithm term, which is the finite part of the loop correction, will not change the poles of the coefficients of the conformal partial wave expansion but will only change the value of the residue at these poles. Therefore, the spectrum keeps invariant under the loop correction as expected. The corrected coefficients can only be solved numerically and we will not show it here.
@@ 6 Multi-collinear limit and the OPE
$\langle 13\rangle=-2\sqrt{\omega_{1}\omega_{3}}\epsilon$ (6.2)
$\langle 23\rangle=-2\sqrt{\omega_{2}\omega_{3}}\delta\epsilon$
$\langle 12\rangle=-2\sqrt{\omega_{1}\omega_{2}}\epsilon(1-\delta)$
From now on we will consider the partial correlator $\mathcal{A}_{n}(\bar{X}X\bar{X}X\cdots)$ without loss of generality  [^5] If one wants to consider the total superamplitudes, just consider to add the integral $\int d^{3}\eta$ with $\eta$ the anti-commuting variable of the cutting leg. The partial fields in $\cdots$ are arbitrary., where $X=X^{4}$ and $\bar{X}=\bar{X}_{4}$. After changing the variables $\omega_{1},\omega_{2},\omega_{3}$ to
@@
We will obtain the following result regardless of the higher $\epsilon$ order terms by cutting a leg and dividing the amplitude into two parts:
[…]
$\frac{(1-s-t)^{\Delta_{3}}}{2(1-\delta)\delta\epsilon}\frac{A_{n-3}\delta^{3}(P)}{st(1-\delta)^{2}+s(1-s-t)+t(1-s-t)\delta^{2}}.$
When we set $\delta\to 0,1,\infty$, we will obtain the result for the consecutive OPE. However, as mentioned before, the collinear limit between 2 momenta is not well-defined, which means that we should keep this $\delta$ to be arbitrary except $0,1,\infty$. Note that this multi-collinear limit cannot be obtained from the 6-point amplitudes since there are other channels besides the channel we considered in the derivation of the multi-collinear limit. Actually, one can show that the multi-collinear limit will lead to an extremely complicated coefficient. Consider the case $\delta\to 0$, the coefficient is
$I_{\Delta_{1}>0}=$ $\frac{-1}{2x_{23}}B(\Delta_{2}-1,\Delta_{3}+1)B(\Delta_{1},\Delta_{2}+\Delta_{3}-1)\times$ (6.5)
${}_{2}F_{1}(1,\Delta_{2}-1,\Delta_{2}+\Delta_{3},\eta)$
when $\Delta_{1}>0$. When $-1<\Delta_{1}<0$, the coefficient becomes a branch cut term:
$I_{-1<\Delta_{1}<0}=$ $\frac{\pi x_{23}^{2\Delta_{1}-1}}{2x_{13}^{2\Delta_{1}}\sin{(\Delta_{1}\pi)}}B(\Delta_{1}+\Delta_{2}-1,\Delta_{1}+\Delta_{3}+1)\times$ (6.6)
${}_{2}F_{1}(\Delta_{1}+1,\Delta_{1}+\Delta_{2}-1,2\Delta_{1}+\Delta_{2}+\Delta_{3},\eta)$
@@
$+\frac{\pi x_{23}^{2\Delta_{1}-1}}{2x_{13}^{2\Delta_{1}}\sin{(\Delta_{1}\pi)}}B(\Delta_{1}+\Delta_{2}-1,\Delta_{1}+\Delta_{3}+1))\bigg{]}X_{\Delta_{1}+\Delta_{2}+\Delta_{3}-1}(x_{3})$
The dominant term of these two terms depends on the value of $\Delta_{1}$, more precisely, the real part of $\Delta_{1}$. The multi-OPE at the higher order can be obtained by expanding some hypergeometric functions and also by considering the descendants of $X$. In multi-OPE, we can find out some information about the OPE. Assuming that $\Delta_{1}$ is in the principal series  [^6] When considering the case $-1<\Delta_{1}<0$, one only needs to consider the following operator, whose descendants also appear in the $X_{\Delta_{2}}(x_{2})\bar{X}_{\Delta_{3}}(x_{3})$ OPE: $\int dx(x-x_{3})^{-2\Delta_{1}}:X\bar{X}:_{\Delta_{2}+\Delta_{3}}(x).$ This OPE is ill-defined since it involves the conformal weight of another operator $\Delta_{1}$. Hence we will not consider this case., we only need to consider $I_{\Delta_{1}>0}$. Since $\delta\to 0$ corresponds to doing the $23$-OPE first, we find that the first term of (6.7) corresponds to the following operator in $X_{\Delta_{2}}(x_{2})\bar{X}_{\Delta_{3}}(x_{3})$ OPE:
$\lim_{\Delta\to 0}\int dx(x-x_{3})^{-2\Delta}:X\bar{X}:_{\Delta_{2}+\Delta_{3}}(x),$ (6.8)
which is a non-local 2-particle operator [21]. And the $1/x_{13}$ pole, which comes from the higher order of the first term, implies the OPE involving a two-particle operator $X\bar{X}\sim:X\bar{X}:$. The corresponding coefficient is
$\frac{-B(\Delta_{2}-1,\Delta_{3}+1)B(\Delta_{1},\Delta_{2}+\Delta_{3}-1)}{2(\Delta_{2}+\Delta_{3})}$ (6.9)
For higher-order of $\delta$, this term will correspond to a certain descendent of $:X\bar{X}:$. Note that from this multi-OPE we find that there will not be an OPE like $X\bar{X}\sim X$ as expected since there will not appear a $1/(x_{13}x_{23})$ pole [^7] It seems that the OPE between 2 outgoing particles and the OPE between 1 incoming particle and 1 outgoing particle are extremely different. This puzzle still needs more clarification in the future..
@@ 7 Discussion
We have demonstrated how to bootstrap the 4-point correlator in the 1D CCFT dual to the ABJM using 3D symmetries as constraints in 1D CFT. We find that the dual inversion property for 4-point correlators, which is also valid for other 1D CFTs, gives a stronger than the crossing symmetry (or cyclic invariant). Meanwhile, from the construction of the dual inversion, we can define some building blocks with given behaviors under the dual inversion and then construct different 4-point correlators with wanting properties under the dual inversion. Furthermore, we calculated the conformal block expansion coefficients to learn more about the spectrum of this CCFT. We then studied the multi-collinear limit of this CCFT from the amplitude to the celestial version and obtained the multi-OPE. In this way, we find a non-local 2-particle operator in the OPE spectrum. This operator, which looks like a certain transformation on the local operator with a kernel, is actually a consequence of the dimensions of fields in 3D. Finally, we try to find out the possible relations between the $\mathcal{N}=4$ SYM and the ABJM with reduced supersymmetry from the CCFT aspect.
@@
There is a relation between $\mathcal{N}=4$ SYM amplitudes and ABJM amplitudes at the tree level from the Grassmannian aspect [34, 35]. Also see [36] for the so-called “orthogonal Grassmannian” which gives the ABJM tree amplitudes. Another interesting question is how to realize it on the celestial sphere. Suppose we believe celestial amplitudes will tell us some information that is hard to find on the amplitude side, it will be interesting to see whether the relation above somehow manifests on the CCFT side. A celestial version of the Grassmannian can be found in [37]. And for celestial $\mathcal{N}=4$ SYM, see [38]. Moreover, at the loop level, the ABJM integrands can be obtained from the $\mathcal{N}=4$ SYM integrands by dimensional reduction [39], can we realize this procedure from the celestial sphere to the celestial circle? This is a similar case to the relation between tree amplitudes as we considered before. Accordingly, we can also ask if there are some relations between these two CCFTs.

### v6

@@ Abstract
In this work, we bootstrap the 4-point correlators on the 1D celestial circle using 3D symmetries in the Aharony-Bergman-Jafferis-Maldacena theory as constraints. We find that the dual inversion property is strong enough to replace the crossing symmetry condition (or cyclic invariant condition) when bootstrapping. We also give some results about the conformal block expansion coefficients which contain the spectrum. Furthermore, we extract the OPE spectrum from the multi-collinear limit since all 3-point ABJM amplitudes vanish. Although we studied a specific theory, the methods used are valid for more general cases.
@@ 1 Introduction and review
Recently, celestial holography, which translates the 4D amplitudes in the bulk to CFT correlators on the celestial sphere that is equivalent to the lightlike infinity of the origin [12, 13], has given us many illustrating results [14, 15, 16, 17, 18]. Inspired by AdS/CFT, celestial holography aims to help us understand quantum gravity. In addition, it also helps us understand other theories from a new aspect. For massless amplitudes, this correspondence can be realized by the Mellin transformation on the magnitude of the external momenta, namely (for 4D amplitudes)
@@
One may ask if we can obtain some new perspectives on the ABJM via CCFT. Furthermore, can we bootstrap the CCFT corresponding to ABJM? In this paper, we will answer these two questions tentatively. We will first show how to bootstrap the 4-point correlators of the CCFT using the 3D symmetry as constraints and then show the conformal block expansion of the 4-point correlator and the loop correction. Finally, we will show the multi-collinear behavior [20, 21] and the in-in (or out-out) OPE of this CCFT. We will see that one can obtain some important information of the CCFT corresponding to ABJM, like the spectrum, from a total CFT perspective.
This paper is organized as follows. In section 2, we will use the celestial version of 3D spin-helicity formalism, which was proposed in [22], to express the 3D symmetries as 1D constraints for tree-level celestial amplitudes [^1] As a supplement, for the celestial version of 4D massive spinor-helicity formalism, see [23].. We will also write down the ansatz for the 4-point tree-level celestial amplitudes. In section 3, we will solve these constraints. In section 4, we will first introduce how the dual inversion acts on the celestial variables for 4-point amplitudes. In section 5, we will investigate the spectrum of this CCFT and the influence of the loop corrections. In section 6, we will consider the multi-collinear limit to obtain the spectrum of the celestial OPE.
@@
$P^{ab}\mathcal{A}_{n}$ $=0\longrightarrow\sum_{i}\epsilon_{i}x_{i}^{j}\mathcal{T}_{i}\mathcal{A}_{n}=0,\text{ for $j=0,1,2$.}$ (2.4)
@@
$Q^{aA}\mathcal{A}_{n}$ $=0\longrightarrow-\sum_{i}\sqrt{-2\epsilon_{i}}\mathcal{T}_{i}^{1/2}\left(\begin{array}[]{c}1\\ x_{i}\end{array}\right)\partial_{\eta_{i}^{A}}\mathcal{A}_{n}=0.$
where the operator $\mathcal{T}_{i}$ can add 1 to the $\Delta_{i}$ in the 1D correlator $\mathcal{A}_{n}$ and the label $j$ is the power of $x_{i}$. In ABJM, besides the usual superconformal symmetry, one can also find the dual superconformal symmetry [10] which will be manifest after using the dual coordinates. Such dual superconformal symmetry can generate a larger algebra together with the original superconformal symmetry, which is a Yangian [24]. This fact is similar to the case in the $\mathcal{N}=4$ SYM, and we only need to consider two level-1 operators in the Yangian. When acting on the 1D correlators, they are equivalent to the dual conformal boost operator and its superpartner. The dual conformal boost operator for ABJM amplitudes reads [10]
@@
    It has the form A_4=g(ϵ_i,Δ_i)f(z,η_i^A)∏_i¡jx_ij^Δ/3-Δ_i-Δ_j with $z=\frac{x_{12}x_{34}}{x_{13}x_{24}}$ the cross-ratio and $\Delta=\sum_{i=1}^{4}\Delta_{i}$.
@@ 3 Four-point correlators from 1D constraints
$\mathcal{A}_{4}[\bar{\Psi}_{1}\Phi_{2}\bar{\Psi}_{3}\Phi_{4}]=\int d\omega\omega^{\sum\Delta_{i}-6}(-\frac{\epsilon_{1}x_{13}x_{14}}{\epsilon_{2}x_{23}x_{24}})^{\Delta_{2}-1}(\frac{\epsilon_{1}x_{12}x_{14}}{\epsilon_{3}x_{23}x_{34}})^{\Delta_{3}-3/2}(-\frac{\epsilon_{1}x_{12}x_{13}}{\epsilon_{4}x_{24}x_{34}})^{\Delta_{4}-2}$ (3.12)
@@ 4 Dual inversion
where
$\Omega_{ij}=\frac{\sqrt{\mathcal{T}_{i}\mathcal{T}_{j}}\bigg(\prod_{k<l}x_{kl}^{\Delta/3-\Delta_{k}-\Delta_{l}}\bigg)}{\prod_{k<l}x_{kl}^{\Delta/3-\Delta_{k}-\Delta_{l}}}$ (4.10)
@@
up to a constant to obtain (4.4). Hence for a 1D CFT, we can impose the constraints (4.4) and also define the dual inversion (LABEL:dual2) and (4.11) for $x_{ij}$ and $\eta_{i}^{A}$ respectively. This constraint, together with the first 3 ansatz in section 2, will determine the 4-point correlator of the 1D CCFT dual to the 3D ABJM theory. Note that the properties (LABEL:dual2) and (4.11) are valid for 4-point correlators in any 1D CFT with the dual conformal symmetry.
@@ 5 Conformal block expansion and loop corrections
Here $\mu$ is the regularization scale and $\zeta(2)=\pi^{2}/6$ is the Riemann zeta function $\zeta(z)$. After acting this operator on the 4-point correlator $\mathcal{A}_{4}$ we will obtain the loop correction. Such a logarithm term, which is the finite part of the loop correction, will not change the poles of the coefficients of the conformal partial wave expansion but will only change the value of the residue at these poles. The corrected coefficients can only be solved numerically and we will not show it here.
@@ 6 Multi-collinear limit and the OPE
[…]
From now on we will consider the partial correlator $\mathcal{A}_{n}(\bar{X}X\bar{X}X\cdots)$ without loss of generality, where $X=X^{4}$ and $\bar{X}=\bar{X}_{4}$. The partial fields in $\cdots$ are arbitrary. We first factorize the $n$-point amplitudes as follows:
$A^{\text{factorized}}_{n}=\int d^{3}\eta A_{4}\frac{1}{p_{123}^{2}}A_{n-3}=\int d^{3}\eta\frac{\delta^{3}(P)}{\langle 12\rangle\langle 23\rangle}\prod_{A=1}^{3}(\sum_{i<j}\langle ij\rangle\eta_{iA}\eta_{jA})\frac{1}{(\langle 12\rangle^{2}+\langle 13\rangle^{2}+\langle 23\rangle^{2})}A_{n-3}$ (6.3)
where $\eta$ is the anti-commuting variable of the cutting leg with momentum $p_{123}$. Focusing on the $\eta_{1}^{3}\eta_{3}^{3}$ term, we have
$A^{\text{factorized}}_{n}(\bar{X}X\bar{X}X\cdots)=-\frac{1}{2\sqrt{\omega_{1}\omega_{2}}(1-\delta)\sqrt{\omega_{2}\omega_{3}}\delta\epsilon}\frac{\omega_{1}\omega_{3}\sqrt{\omega_{1}\omega_{3}}}{\omega_{1}\omega_{2}(1-\delta)^{2}+\omega_{1}\omega_{3}+\omega_{2}\omega_{3}\delta^{2}}A_{n-3}$ (6.4)
After changing the variables $\omega_{1},\omega_{2},\omega_{3}$ to
@@
We will obtain the following celestial result regardless of the higher $\epsilon$ order terms [^5] We have also set $\epsilon=0$ in the delta function.:
$\mathcal{A}^{\text{factorized}}_{n}=$ $-\int_{0}^{\infty}d\omega\int_{0}^{1}ds\int_{0}^{1-s}dt\omega^{\Delta_{1}+\Delta_{2}+\Delta_{3}-2}s^{\Delta_{1}}t^{\Delta_{2}-2}$ (6.6)
$\times\frac{(1-s-t)^{\Delta_{3}}}{2(1-\delta)\delta\epsilon}\frac{A_{n-3}\delta^{3}(P)}{st(1-\delta)^{2}+s(1-s-t)+t(1-s-t)\delta^{2}}=I\mathcal{A}_{n-3}.$
with
$I=-\int_{0}^{1}ds\int_{0}^{1-s}dt\frac{(1-s-t)^{\Delta_{3}}}{2(1-\delta)\delta\epsilon}\frac{s^{\Delta_{1}}t^{\Delta_{2}-2}}{st(1-\delta)^{2}+s(1-s-t)+t(1-s-t)\delta^{2}}$ (6.7)
This formula gives the following leading multi-OPE
$\bar{X}_{\Delta_{1}}X_{\Delta_{2}}\bar{X}_{\Delta_{3}}\sim\bar{X}_{\Delta_{1}+\Delta_{2}+\Delta_{3}-1}$ (6.8)
with the multi-OPE prefactor $I$ (including the pole factor and the multi-OPE coefficient). When we set $\delta\to 0,1,\infty$, we will obtain the result for the consecutive OPE. However, as mentioned before, the collinear limit between 2 momenta is not well-defined, which means that we should keep this $\delta$ to be arbitrary except $0,1,\infty$. Note that this multi-collinear limit cannot be obtained from the 6-point amplitudes directly since there are other channels besides the channel we considered in the derivation of the multi-collinear limit. Consider the case $\delta\to 0$, the prefactor is
$I_{\Delta_{1}>0}=$ $\frac{-1}{2x_{23}}B(\Delta_{2}-1,\Delta_{3}+1)B(\Delta_{1},\Delta_{2}+\Delta_{3}-1)\,{}_{2}F_{1}(1,\Delta_{2}-1,\Delta_{2}+\Delta_{3},2\delta)$ (6.9)
when $\Delta_{1}>0$. Here we only ignore the $\delta^{2}$ terms in the denominator. When $1<\Delta_{1}<0$, the prefactor becomes a branch cut term:
$I_{-1<\Delta_{1}<0}=$ $\frac{\pi x_{23}^{2\Delta_{1}-1}}{2x_{13}^{2\Delta_{1}}\sin{(\Delta_{1}\pi)}}B(\Delta_{1}+\Delta_{2}-1,\Delta_{1}+\Delta_{3}+1)\,{}_{2}F_{1}(\Delta_{1}+1,\Delta_{1}+\Delta_{2}-1,2\Delta_{1}+\Delta_{2}+\Delta_{3},2\delta)$ (6.10)
@@
$+\frac{\pi x_{23}^{2\Delta_{1}-1}}{2x_{13}^{2\Delta_{1}}\sin{(\Delta_{1}\pi)}}B(\Delta_{1}+\Delta_{2}-1,\Delta_{1}+\Delta_{3}+1))\bigg]\bar{X}_{\Delta_{1}+\Delta_{2}+\Delta_{3}-1}(x_{3})$
The dominant term of these two terms depends on the value of $\Delta_{1}$, more precisely, the real part of $\Delta_{1}$. The multi-OPE at the higher order can be obtained by expanding some hypergeometric functions and also by considering the descendants of $X$. In multi-OPE, we can find out some information about the OPE. Assuming that $\Delta_{1}$ is in the principal series [^6] When considering the case $1<\Delta_{1}<0$, one only needs to consider the singularity $x_{23}^{2\Delta_{1}-1}$, which depends on $\Delta_{1}$. Hence, this OPE is ill-defined., we only need to consider $I_{\Delta_{1}>0}$. Since $\delta\to 0$ corresponds to doing the $23$-OPE first, we find that the first term of (6.11) corresponds to the following operator in $X_{\Delta_{2}}(x_{2})\bar{X}_{\Delta_{3}}(x_{3})$ OPE by the conformal weight analysis:
$X_{\Delta_{2}}(x_{2})\bar{X}_{\Delta_{3}}(x_{3})\sim\frac{C}{x_{23}}\mathcal{O}_{\Delta_{2}+\Delta_{3}-1}(x_{3})$ (6.12)
with a constant $C$. However, all 3-point amplitudes in ABJM vanish; hence, the operator $\mathcal{O}$ must be a 2-particle operator [21] in this leading OPE. Note that we also have 4-point correlators with 2 bosonic fields and 2 fermionic fields. The only choice for such a primary operator is
$\mathcal{O}_{\Delta_{2}+\Delta_{3}-1}\sim\ \text{combination of some}\ :\bar{\psi}\psi:_{\Delta_{2}+\Delta_{3}-1}$ (6.13)
where $\psi^{\mathsf{A}}$ are the complex fermionic fields in the ABJM. This observation matches the Lagrangian of ABJM.
By expanding the hypergeometric function w.r.t $\delta$, we find that the poles in the multi-OPE are $x_{23}^{n-1}/x_{13}^{n}$ with $n=0,1,\cdots$ the order of $\delta$. The $1/x_{13}$ pole, which comes from the higher order of the first term, implies the OPE involving a 2-particle operator $X\bar{X}\sim:X\bar{X}:$ and $X\bar{X}\sim\partial_{x}\mathcal{O}$:
$\bar{X}_{\Delta_{1}}(x_{1})\big[C\partial\mathcal{O}(x_{3})+:X_{\Delta_{2}}\bar{X}_{\Delta_{3}}:(x_{3})\big]\sim-\frac{1}{x_{13}}B(\Delta_{2},\Delta_{3}+1)B(\Delta_{1},\Delta_{2}+\Delta_{3}-1)X_{\Delta_{1}+\Delta_{2}+\Delta_{3}-1}(x_{3})$ (6.14)
For higher $\delta$ orders, these poles will correspond to some descendants of $X\bar{X}:$ and $\mathcal{O}$. In this case, the $\delta^{2}$ terms in (6.7) cannot be ignored, and we need to expand the $\delta^{2}$ terms in it. However, it is still unknown how to obtain the specific OPE coefficients of the $X\bar{X}$ OPE as there are two types of operators :$X$::$X$: and $\mathcal{O}$ and their corresponding coefficients mix. We leave this to future work.
It is important to note that from this multi-OPE we find that there will not be an OPE like $X\bar{X}\sim X$ as expected, since there will not appear a $1/(x_{13}x_{23})$ pole. Then, from the multi-OPE perspective, we find the interaction information from the spectrum.
@@ 7 Discussion
We have demonstrated how to bootstrap the 4-point correlator in the 1D CCFT dual to the ABJM using 3D symmetries as constraints in 1D CFTs. We find that the dual inversion property for 4-point correlators, which is also valid for other 1D CFTs with dual conformal symmetries, gives a stronger constraint than the crossing symmetry (or cyclic invariant). Meanwhile, from the construction of the dual inversion, we can define some building blocks with given behaviors under the dual inversion and then construct different 4-point correlators with desired properties under the dual inversion. Furthermore, we calculated the conformal block expansion coefficients to learn more about the spectrum of this CCFT. We then studied the multi-collinear limit of this CCFT from the amplitude to the celestial version and obtained the multi-OPE. In this way, we find a special 2-particle operator in the OPE spectrum. Note that we only obtain the spectrum of OPE from the multi-OPE; the specific coefficients of each operator are still unknown since different operators will contribute to the same term in the multi-OPE. How to obtain the whole $X\bar{X}$ OPE will be left to future work.
@@
There is a relation between $\mathcal{N}=4$ SYM amplitudes and ABJM amplitudes at the tree level from the Grassmannian aspect [34, 35]. Also see [36] for the so-called “orthogonal Grassmannian” which gives the ABJM tree amplitudes. Another interesting question is how to realize it on the celestial sphere. Suppose we believe celestial amplitudes will tell us some information that is hard to find on the amplitude side, it will be interesting to see whether the relation above somehow manifests on the CCFT side. A celestial version of the Grassmannian can be found in [37]. And for celestial $\mathcal{N}=4$ SYM, see [38, 39]. Moreover, at the loop level, the ABJM integrands can be obtained from the $\mathcal{N}=4$ SYM integrands by dimensional reduction [40], can we realize this procedure from the celestial sphere to the celestial circle? This is a similar case to the relation between tree amplitudes as we considered before. Accordingly, we can also ask if there are some relations between these two CCFTs.